An electron (charge= -1.6 x 10^-19 C) is moving at 3 x 10^5 m/s in the positive x direction. A magnetic field of 0.8 T is in the positive z direction. What are the magnitude and direction on the magnetic force on the electron?
Answer #1

Force, F=q V x B

= -1.6*10-19 (3*105i x0.8k)

= -3.84*10-14 (-j)

= 3.84*10-14j

The magmitude of force is 3.84*10-14 N and direction is + y.